# Understanding Oversmoothing in GNNs as Consensus in Opinion Dynamics

### v1

@@ 5.3 Designing an expressive BIMP model
The critical point of the attention-opinion bifurcation diagram occurs when the bias parameter $\mathbf{B}$ is equal to zero and agent opinions are at equilibrium.
Following (Leonard et al., 2024), we use a linear analysis to determine the critical point of the attention-opinion bifurcation diagram. The linearization of the nonlinear opinion dynamics model (Equation (24)) is given by $\dot{\boldsymbol{\omega}}=J(\mathbf{x}_{e})\,\boldsymbol{\omega}$, where $J(\mathbf{x}_{e})$ is the Jacobian evaluated at the equilibrium $\mathbf{x}_{e}$, and $\boldsymbol{\omega}=\mathbf{x}-\mathbf{x}_{e}$. At the equilibrium $\mathbf{x}_{e}=\mathbf{0}$, the Jacobian is given by
$\mathbf{J}=(-d+u\tilde{\alpha})\mathbf{I}+u\tilde{\mathbf{A}}.$ (27)
$\lambda^{\prime}=-d+u\tilde{\alpha}+u\lambda^{\tilde{a}}.$ (28)
The critical point of the attention-opinion bifurcation diagram occurs when the dominant eigenvalue of the Jacobian is zero. Solving for the critical value of the attention yields $u=\nicefrac{{d}}{{\tilde{\alpha}+\lambda^{\tilde{a}}_{\max}}}$. ∎
To obtain a value for the attention, we must know the value for $d$, $\alpha$, and $\lambda^{\tilde{a}}_{\max}$. While we treat $d$ and $\alpha$ as hyperparameters of our model, $\lambda^{\tilde{a}}_{\max}$ is determined by the structure of our effective adjacency matrix.
Proposition 5.5.
The dominant eigenvalue $\lambda^{\tilde{a}}_{\max}$ of $\tilde{\mathbf{A}}$ is equal to four.
Proof.
Since both the communication matrix $\mathbf{A}^{\mathrm{a}}$ and the belief matrix $\mathbf{A}^{\mathrm{o}}$ are right stochastic, their dominant eigenvalues are equal to one (i.e., $\lambda^{\mathrm{a}}_{\max}=\lambda^{\mathrm{o}}_{\max}=1$). Since $\tilde{\mathbf{A}}=(\mathbf{A}^{\mathrm{o}}+I)\otimes(\mathbf{A}^{\mathrm{a}}+I)$, its dominant eigenvalue is equal to $\lambda^{\tilde{a}}_{\max}=(\lambda^{\mathrm{a}}_{\max}+1)(\lambda^{\mathrm{o}}_{\max}+1)=4$. ∎
Leveraging Proposition 5.4 and 5.5, we have that $u=\nicefrac{{d}}{{\alpha+3}}$. Notably, the attention $u$ is independent of the learned values of $\mathbf{A}^{\mathrm{a}}$ and $\mathbf{A}^{\mathrm{o}}$. This significantly reduces computational complexity since we do not need to recompute $\lambda^{\tilde{a}}_{\max}$ and $u$ each time $\mathbf{A}^{\mathrm{a}}$ or $\mathbf{A}^{\mathrm{o}}$ is updated.
BIMP equilibrium behavior.
With our choice of effective adjacency matrix $\tilde{\mathbf{A}}$ and attention parameter $u$, BIMP converges to equilibrium when the input parameter $\mathbf{B}$ is constant.
Proposition 5.6.
A BIMP model with time-independent bias parameter $\mathbf{B}$ converges to an equilibrium.
We prove Proposition 5.6 in Appendix A.2 by showing that the nonlinear opinion dynamics are monotone and bounded. Consequently, there are no dynamical behaviors other than convergence to an equilibrium. Since all trajectories tend toward the equilibrium solution, analyzing the equilibrium behavior is sufficient to understand the dynamics of BIMP.
The bias parameter
We can show that if the input parameter is chosen to be nonzero, BIMP can achieve dissensus (i.e., will not exhibit oversmoothing). We first explore the behavior of BIMP when the input parameter $\mathbf{B}$ is equal to zero.
Proposition 5.7.
When the bias parameter $\mathbf{B}$ is equal to zero, the dynamics of BIMP can be approximated by the dynamics of the reduced one-dimensional dynamical equation.
$\dot{y}=-d\,y+\operatorname{tanh}\left[u(\tilde{\alpha}+4)y\right],$ (29)
where $y=\langle\mathbf{x},\mathbf{w}_{\max}\rangle\in\mathbb{R}$, and $\mathbf{w}_{\max}$ is the left dominant eigenvector of $\tilde{\mathbf{A}}$.
Following (Leonard et al., 2024), we use the Lyapunov–Schmidt reduction (Golubitsky & Schaeffer, 1985) to determine the reduced model. We decompose our BIMP’s dynamics into a critical subspace and a regular subspace. The critical subspace corresponds to the zero eigenvalue of linearized system, which governs the long-term behavior (i.e., equilibrium). In contrast, the regular subspace corresponds to the negative eigenvalues of the linearized system, which decays quickly and have a limited contribution to the long-term dynamics. It is therefore sufficient to focus on the critical subspace to understand the dynamics of the equilibrium as the regular subspace decays quickly. We derive Equation (47) in Appendix A.3.
Proposition 5.8.
BIMP exhibits oversmoothing when the bias parameter $\mathbf{B}$ is equal to zero.
Proof.
For $\mathbf{x}$ in the neighborhood of the equilibrium $\mathbf{x}=\mathbf{0}$, the right-hand side of Equation (47) is isomorphic to $(u(\tilde{\alpha}+4)-d)y-u(\tilde{\alpha}+4)y^{3}$, and at the critical point $u=\nicefrac{{d}}{{\alpha+3}}$ Equation (47) has unique equilibrium $y=0$. This corresponds to an equilibrium solution of $\mathbf{x}=\mathbf{0}$ in the original system (Equation (24)) which means that $\mathbf{X}=\mathbf{0}$ and all agents form neutral opinions for all options. Since the opinions of all agents have converged, the system has reached consensus (i.e., exhibits oversmoothing). ∎
Proposition 5.9.
BIMP will not exhibit oversmoothing when the bias parameter $\mathbf{B}$ is time-independent with unique entries.
A distinct and time-independent $\mathbf{B}$ serves as a constant forcing term that applies a continuous influence on each opinion $\mathbf{x}_{i}$, biasing the evolution of $\mathbf{X}$ toward $\mathbf{B}$. Consequently, the system equilibrium is shifted and opinions reach dissensus. We prove Proposition 5.9 in Appendix A.4.
To ensure $\mathbf{B}$ satisfies the conditions of Proposition 5.9, we we set $\mathbf{B}$ to the initial opinion state, i.e., $\mathbf{B}=\mathbf{X}(0)$. A similar strategy is used GRAND++ (Thorpe et al., 2022) and KuramotoGNN (Nguyen et al., 2024).
Figure 1: Robustness to oversmoothing. Classification accuracy for increasing network depth is shown for the Cora, Citeseer, and Pubmed datasets (left-right). In all cases, BIMP shows negligible changes in accuracy while competitive baselines exhibit significantly reduced accuracy.
Table 2: Classification accuracy. Classification accuracy on the Cora, Citeseer, Pubmed, CoauthorCS, Computers and Photo datasets is reported. Our BIMP model outperforms competitive baselines on all datasets.
Dataset BIMP (ours) $-\ell$ $-\ell$ KuramotoGNN GraphCON GAT GCN GraphSAGE
Cora 83.19$\pm$1.13 82.20$\pm$1.45 82.83$\pm$1.31 81.16$\pm$1.61 82.80$\pm$1.34 79.76$\pm$1.50 80.76$\pm$2.04 79.37$\pm$1.70
Citeseer 71.09$\pm$1.40 69.89$\pm$1.48 70.26$\pm$1.46 70.40$\pm$1.02 69.60$\pm$1.16 67.70$\pm$1.63 67.54$\pm$1.98 67.31$\pm$1.63
Pubmed 80.16$\pm$2.03 78.19$\pm$1.88 78.89$\pm$1.96 78.69$\pm$1.91 78.85$\pm$1.53 76.88$\pm$2.08 77.04$\pm$1.78 75.52$\pm$2.19
CoauthorCS 92.48$\pm$0.26 90.23$\pm$0.91 90.10$\pm$0.78 91.05$\pm$0.56 90.30$\pm$0.74 89.51$\pm$0.54 90.98$\pm$0.42 90.62$\pm$0.42
Computers 84.73$\pm$0.61 82.93$\pm$0.56 82.79$\pm$0.54 80.06$\pm$1.60 82.76$\pm$0.58 81.73$\pm$1.89 82.02$\pm$1.87 76.42$\pm$7.60
Photo 92.90$\pm$0.44 91.93$\pm$0.39 91.51$\pm$0.41 92.77$\pm$0.42 91.78$\pm$0.50 89.12$\pm$1.60 90.37$\pm$1.38 88.71$\pm$2.68
5.4 BIMP training behavior
We can show that even for very deep networks, BIMP gradients will not explode or exponentially vanish during training.
Proposition 5.10.
BIMP gradients are upper-bounded when the step-size $\Delta t\ll 1$ and damping term $d<\nicefrac{{1}}{{\Delta t}}$.
Proposition 5.11.
Therefore, BIMP maintains well behaved gradients even for very deep architectures. We prove Proposition 5.10 and 5.11 in Appendix A.5.
6 Experiments
In this section, we highlight our model’s improved robustness to oversmoothing compared to competitive baselines, strong performance on graph classification tasks, and resilience to adversarial attack.
6.1 Robustness to oversmoothing
We measure oversmoothing in GCN (Kipf & Welling, 2016), GAT (Veličković et al., 2017), $-\ell$ (Chamberlain et al., 2021b), $-\ell$ (Thorpe et al., 2022), GraphCON-Tran (Rusch et al., 2022) and our BIMP model across network depths using the Dirichlet energy (see Equation (2)), and plot the results in Figure 2. The Dirichlet energy of our BIMP model is stable out to 1000 layers while the energy of other models decreases (i.e., approaches oversmoothing). Additional experimental details are provided in Appendix C.2.1.
To understand the impact of oversmoothing on classification accuracy, we compare the performance of our BIMP model, $-\ell$ , $-\ell$ and GraphCON-Tran on the Cora (McCallum et al., 2000), Citeseer (Sen et al., 2008) and Pubmed (Namata et al., 2012) datasets. Each model is trained over a range of network depths and fixed hyperparameters; results are reported in Figure 1. Our BIMP model has comparable performance with baseline models when networks depths are small, but increasingly better performance compared to baseline models as network depths increase. Additional experimental details are provided in Appendix C.2.2.
Figure 2: Dirichlet energy. We compare the Dirchelet energy of node features over a range of network depths. The Dirchelet energy of BIMP models remains stable even at very deep layers, consistently outperforming baseline models.
6.2 Improved classification accuracy
We demonstrate improved classification performance compared to $-\ell$, GRAND+++++ $-\ell$, KuramotoGNN, GraphCON, GAT, GCN, and GraphSAGE on the Cora, Citeseer, Pubmed, Computers (Shchur et al., 2018), CoauthorCS (Shchur et al., 2018), and Photo (Shchur et al., 2018) datasets. Classification accuracies are reported in Table 2, where BIMP consistently outperforms baseline models. We also observe the improved performance on heterophilic datasets reported in Table 7 in Appendix D.2.
6.3 Resilience to adversarial attack
We demonstrate improved resilience to adversarial attack compared to $-\ell$ under random and PGD (Mkadry et al., 2017) attack methods on the Cora, Citeseer and Pubmed datasets. We use an untargeted graph modification attack, perturbingnode features at inference. In the random attack, we sample the perturbation from $\mathcal{N}(0,1)$, and in the PGD attack, we choose $\epsilon=0.03,\alpha=0.01$. Classification accuracies are reported in Table 3, and show that our BIMP model significantly outperforms the baseline model.
6.4 Computational complexity
We compare the space complexity of our BIMP model with that of $-\ell$. The complexity our model is $\mathcal{O}(mN_{o}+N_{o}^{2}N_{\mathrm{a}})$, where $m$ is the number of edges, $N_{\mathrm{a}}$ is the number of nodes, and $N_{\mathrm{o}}$ is the feature dimension. The space complexity of our model is comparable to $-\ell$, which has complexity $\mathcal{O}(|\mathcal{E}|N_{o})$).
We also compute the time complexity of our model. We compare the average run time of our model against baseline models in Table 6 of Appendix C.4. Despite incorporating nonlinearity and belief matrix $\mathbf{A}^{\mathrm{o}}$, BIMP has comparable run time performance to linear baselines models. Additional experimental details are provided in Appendix C.4.
Table 3: Robustness to adversarial attack. Classification accuracy under different types of adversarial attacks are reported. BIMP shows significantly improved resilience to adversarial attacks compared to linear baseline models.
Dataset Attack method BIMP (ours) $-\ell$
Cora Clean 82.4$\pm$2.1 82.1$\pm$1.4
Random 70.8$\pm$1.6 61.6$\pm$2.3
PGD 56.4$\pm$5.4 38.4$\pm$4.9
Citeseer Clean 70.4$\pm$1.6 70.0$\pm$1.7
Random 57.8$\pm$1.2 42.5$\pm$2.2
PGD 57.4$\pm$3.1 39.6$\pm$3.7
Pubmed Clean 79.7$\pm$1.9 78.3$\pm$2.1
Random 61.2$\pm$6.7 40.8$\pm$2.5
PGD 40.1$\pm$7.8 37.6$\pm$5.6
7 Conclusion
In this paper, we propose an analogy between GNNs and opinion dynamics models highlighting the equivalence between the oversmoothing condition in GNNs and consensus condition in opinion dynamics. Through our analogy, we show that recent continuous-depth GNNs are equivalent to linear opinion dynamics models, which are proven to converges to consensus. To address this issue, we introduce a novel class of continuous-depth GNNs called Behavior-inspired message passing (BIMP) neural networks which leverage the nonlinear opinion dynamics inductive bias. The nonlinear saturation function and distinct bias in BIMP results in greater expressivity and guaranteed dissensus. Experiments against competitive baselines illustrate our model’s robustness to network depth, improved expressivity, and resiliance to adversarial attack.
Limitation and Future work
With large step sizes, the damping term $d$ leads to training instability, limiting training acceleration with large time scales. Future work will develop deeper insights into the effect of the nonlinearity in BIMP on its resilience to adversarial attacks.
Impact statement
We present a novel continuous-depth graph neural networks to tackle the oversmoothing issue and enable us to build a deep model with enhanced expressive capacity. The improved expressivity can lead to some breakthroughs in various domains such as drug discovery and protein prediction, where we don’t anticipate the negative societal impacts.
Appendix A Proof of Propositions
A.1 Proposition 4.2: Solution for the system with source term
As $t\to\infty$ the dynamics of $-\ell$ are approximately linear in $t$.
Proof.
Consider the dynamics of $-\ell$ given by
@@ Appendix A Proof of Propositions
where $\mathbf{L=I-A}$ is the generalized graph Laplacian, $\mathbf{b}$ is the fixed source term. Defining the right eigenvector matrix $\mathbf{T}=[\mathbf{v}_{1},\mathbf{v}_{2},...,\mathbf{v}_{N_{a}}]$ and left eigenvector matrix $\mathbf{T}^{-1}=[\mathbf{w}_{1},\mathbf{w}_{2},...,\mathbf{w}_{N_{a}}]^{\top}$ for the generalized Laplacian $\mathbf{L}$, perform change of coordinates $\mathbf{Y}=[\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{N_{\mathrm{a}}}]^{\top}=\mathbf{T}^{-1}\mathbf{X}$ such that
@@
$\mathbf{T}\dot{\mathbf{Y}}$ $=-\mathbf{T\Lambda T}^{-1}\mathbf{T}\mathbf{Y}+\mathbf{B},$ (32)
$=-\mathbf{T\Lambda}\mathbf{Y}+\mathbf{B}.$ (33)
@@
where $\lambda_{i}$ is the $i$-th eigenvalue of $\mathbf{\Lambda}$. Consider the case where $\lambda_{i}>0$, the solution to the ODE becomes
@@
where $\lambda_{i}$ and $\mathbf{v}_{i}$ denote the positive eigenvalues and corresponding eigenvectors of $\mathbf{L}$; $\mathbf{b}_{i}=\mathbf{w}_{i}^{\top}\mathbf{B}$, where $\mathbf{w}_{i}$ are the left eigenvectors of $\mathbf{L}$. Particularly, $\mathbf{v}_{0}$ is an all-ones vector, i.e., $\mathbf{v}_{0}=[1,1,..,1]^{\top}$, and $\mathbf{b}_{0}=\mathbf{w}_{0}^{\top}\mathbf{B}$. $\mathbf{c}_{0}$, $\mathbf{c}_{i}$ are some constant vectors. As $t\rightarrow\infty$, the exponential terms decay and Equation (38) tends towards a linear system dominated by $\mathbf{v}_{0}$. ∎
A.2 Proposition 5.6: BIMP converges to equilibrium

### v2

@@ A.1 Oversmoothing and Opinion Consensus
In $-\ell$, the layer-wise transformations is of the form,
@@
where $\mathbf{B}$ is the fixed source term that depends on the initial state $\mathbf{X}(0)$. Defining the right eigenvector matrix $\mathbf{T}=[\mathbf{v}_{1},\mathbf{v}_{2},...,\mathbf{v}_{N_{a}}]$ and left eigenvector matrix $\mathbf{T}^{-1}=[\mathbf{w}_{1},\mathbf{w}_{2},...,\mathbf{w}_{N_{a}}]^{\top}$ for the graph Laplacian $\mathbf{L}$, we perform change of coordinates $\mathbf{Y}=[\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{N_{\mathrm{a}}}]^{\top}=\mathbf{T}^{-1}\mathbf{X}$ such that
@@
$\mathbf{T}\dot{\mathbf{Y}}$ $=-\mathbf{T\Lambda T}^{-1}\mathbf{T}\mathbf{Y}+\mathbf{B}=-\mathbf{T\Lambda}\mathbf{Y}+\mathbf{B}.$ (21)
@@
where $\lambda_{i}$ is the $i$-th eigenvalue of $\mathbf{\Lambda}$. Consider the case of $\lambda_{i}>0$, the solution to the ODE becomes
@@
where $\lambda_{i}$ and $\mathbf{v}_{i}$ denote the positive eigenvalues and corresponding Eigenvectors of $\mathbf{L}$; $\mathbf{b}_{i}=\mathbf{w}_{i}^{\top}\mathbf{B}$, where $\mathbf{w}_{i}$ are the left eigenvectors of $\mathbf{L}$. Particularly, $\mathbf{v}_{0}$ is an all-ones vector, i.e., $\mathbf{v}_{0}=[1,1,..,1]^{\top}$, $\mathbf{b}_{0}=\mathbf{w}_{0}^{\top}\mathbf{B}$, and $\mathbf{c}_{0}$, $\mathbf{c}_{i}$ are constant vectors. As $t\rightarrow\infty$, the exponential terms $\mathbf{c}_{i}^{\top}e^{-\lambda_{i}t}$ decays to zero and Equation (26) tends towards a linear system dominated by $\mathbf{v}_{0}$. Moreover, for large timescales, the difference in node features is relatively small resulting in reduced discriminability (another form of oversmoothing), and poorer network performance. ∎
Lemma A.2 (Oversmoothing in GREAD-F and GREAD-FB*).
GREAD-F and GREAD-FB exhibits oversmoothing.
Proof.
GREAD-F. In GREAD-F, the layer-wise transformations is of the form,
$\dot{\mathbf{X}}=-\mathbf{L}\mathbf{X}+\mathbf{X}\odot(1-\mathbf{X}),$ (27)
where $\odot$ denotes the Hadamard product. For Laplacian $\mathbf{L}$, there exist a constant $C>0$ such that
$|[\mathbf{LX}]_{i}|\leq C|\mathbf{X}_{i}|,$ (28)
where $C$ depends on the maximum degree and largest edge weights and $\mathbf{X}_{i}$ denotes the $i$-th row of $\mathbf{X}$. Equation (27) can therefore be rewritten as
$\dot{\mathbf{X}_{i}}=-[\mathbf{LX}]_{i}+\mathbf{X}_{i}(1-\mathbf{X}_{i})\leq C|\mathbf{X}_{i}|+\mathbf{X}_{i}-\mathbf{X}^{2}_{i}.$ (29)
Notably, when $\mathbf{X}_{i}<-C$, the RHS of Equation (29) is strictly smaller than zero. Since the dominant term $\mathbf{X}_{i}^{2}$ grows quadratically, the solution diverges to infinity
$\lim_{t\to\infty}\mathbf{X}_{i}(t)=-\infty\quad\text{for all }i.$ (30)
Therefore, as long as the maximum of $\mathbf{X}$ is smaller than some negative threshold $-C$, the entire system becomes monotonically decreasing, eliminating the possibility of equilibrium or steady-state convergence. While node values may diverge at different rates, the components with the fastest decay rate dominates the overall behavior. The remaining components which decays more slowly becomes negligible in relative magnitude. Therefore, all node features collapse and oversmoothing occurs.
GREAD-FB. In GREAD-FB, the layer-wise transformations is of the form,
$\dot{\mathbf{X}}=-\alpha\mathbf{L}\mathbf{X}+\beta(\mathbf{LX}+\mathbf{X}),$ (31)
where $\alpha,\beta$ are trainable parameters to (de-)emphasize each term. As the external input is a linear transformation, the dynamics can be rewritten as
$\dot{\mathbf{X}}=\underbrace{\biggl{(}(\beta-\alpha)\mathbf{L}+\beta\mathbf{I}\biggr{)}}_{\tilde{\mathbf{L}}}\mathbf{X}.$ (32)
Given the property of linear opinion dynamics, when $\alpha>\beta>0$, the eigenvalue of $\tilde{\mathbf{L}}$ exists in range $[2\Delta(\beta-\alpha),\beta]$, where $\Delta$ is the maximum graph degree. In particular, the smallest eigenvalue of $\mathbf{L}$, which is zero, maps to the largest eigenvalue of $\lambda_{\tilde{\mathbf{L}}}=\beta$. This guarantees a global stable equilibrium cannot exist as at least one mode grows exponentially with rate $\beta$.
Meanwhile, as $t\to\infty$, the remaining components associated with smaller eigenvalues either increases slowly or decays to zero (depends on the sign of the eigenvalue), and thus becoming negligible in relative magnitude. Consequently, all node features are asymptotically dominated by the leading eigenvector, which is $[1,\cdots,1]^{\top}$. This leads to a collapse in feature diversity and oversmoothing occurs. ∎
A.2 Behavior-Inspired Message Passing Neural Network
A.2.1 Lemma 5.2
The general form of nonlinear opinion dynamics (Equation (11)), $\dot{\mathbf{X}}(t)=-d\mathbf{X}(t)+\tanh\left[u\left(\alpha\mathbf{X}(t)+\mathbf{A}^{\mathrm{a}}\mathbf{X}(t)+\mathbf{X}(t){\mathbf{A}^{\mathrm{o}}}^{\top}+\mathbf{A}^{\mathrm{a}}\mathbf{X}(t){\mathbf{A}^{\mathrm{o}}}^{\top}\right)\right]+\mathbf{X}(0),$ can be written, $\dot{\mathbf{x}}=-d\mathbf{x}+\tanh\left[u\biggl{(}(\alpha-1)\mathbf{x}+\tilde{\mathbf{A}}\mathbf{x}\biggr{)}\right]+\mathbf{b},$ (33) where $\tilde{\mathbf{A}}$ is the effective adjacency matrix, $\mathbf{x}=\operatorname{vec}(\mathbf{X})$ and $\mathbf{b}=\operatorname{vec}(\mathbf{B})$.
Proof.
We first rewrite Equation (11) as
$\dot{\mathbf{X}}=-d\mathbf{X}+\tanh\Big{[}u\Big{(}(\alpha-1)\mathbf{X}+(\mathbf{A}^{\mathrm{a}}+\mathbf{I})\mathbf{X}({\mathbf{A}^{\mathrm{o}}}^{\top}+\mathbf{I})\Big{)}\Big{]}+\mathbf{B}.$ (34)
From here, we can write the matrix product $\mathbf{ABC}$ with $\mathbf{A}\in\mathbb{R}^{m\times m}$, $\mathbf{B}\in\mathbb{R}^{m\times n}$ and $\mathbf{C}\in\mathbb{R}^{n\times n}$, as $\mathbf{ABC}=\operatorname{vec}^{-1}\left[\left(\mathbf{C}^{\top}\otimes\mathbf{A}\right)\operatorname{vec}(\mathbf{B})\right]$, where $\operatorname{vec}$ denotes the vectorization operator [91]. This yields the following form
$\dot{\mathbf{X}}=-d\mathbf{X}+\tanh\left[u\left((\alpha-1)\mathbf{X}+\operatorname{vec^{-1}}(\tilde{\mathbf{A}}\operatorname{vec}(\mathbf{X})\right)\right]+\mathbf{B},$ (35)
where $\tilde{\mathbf{A}}=(\mathbf{A}^{\mathrm{o}}+\mathbf{I})\otimes(\mathbf{A}^{\mathrm{a}}+\mathbf{I})$ follows from Definition 5.1. Vectorizing both sides of yields
$\dot{\mathbf{x}}=-d\mathbf{x}+\tanh\left[u\left((\alpha-1)\mathbf{x}+\tilde{\mathbf{A}}\mathbf{x}\right)\right]+\mathbf{b},$ (36)
where $\mathbf{x}=\operatorname{vec}(\mathbf{X})$, and $\mathbf{b}=\operatorname{vec}(\mathbf{B})$. We obtain the general nonlinear opinion dynamics in the form of Equation (12). By vectorizing $\mathbf{X}$, each agent opinion on an option is treated as an individual agent-like entity, thereby reducing the original dynamics to the form where options are uncorrelated. ∎
A.2.2 Lemma 5.3
The leading eigenvalue of the effective adjacency matrix $\tilde{\mathbf{A}}$ is $\lambda^{\tilde{a}}_{\max}=4$.
Proof.
Since both the communication adjacency matrix $\mathbf{A}^{\mathrm{a}}$ and the belief adjacency matrix $\mathbf{A}^{\mathrm{o}}$ are right stochastic, their leading eigenvalues are equal to one (i.e., $\lambda^{\mathrm{a}}_{\max}=\lambda^{\mathrm{o}}_{\max}=1$). Since $\tilde{\mathbf{A}}=(\mathbf{A}^{\mathrm{o}}+\mathbf{I})\otimes(\mathbf{A}^{\mathrm{a}}+\mathbf{I})$, its leading eigenvalue is equal to $\lambda^{\tilde{a}}_{\max}=(\lambda^{\mathrm{a}}_{\max}+1)(\lambda^{\mathrm{o}}_{\max}+1)=4$. ∎
A.2.3 Lemma 5.4: Bifurcation point $u^{*}$
The bifurcation point $u^{*}$ of the attention-opinion bifurcation diagram is equal to $\nicefrac{{d}}{{\left(\alpha-1+\lambda^{\tilde{a}}_{\max}\right)}}$.
Proof.
The bifurcation point of the attention-opinion bifurcation diagram, i.e., the point where the number and/or stability of the solutions change, occurs when the input parameter $\mathbf{B}$ is equal to zero.
Following [35], we use a linear analysis to determine the bifurcation point of the attention-opinion bifurcation diagram. The linearization of the nonlinear opinion dynamics model (Equation (36)) is given by $\dot{\boldsymbol{\omega}}=J(\mathbf{x}_{e})\,\boldsymbol{\omega}$, where $J(\mathbf{x}_{e})$ is the Jacobian evaluated at the equilibrium $\mathbf{x}_{e}$, and $\boldsymbol{\omega}=\mathbf{x}-\mathbf{x}_{e}$. At the equilibrium $\mathbf{x}_{e}=\mathbf{0}$, the Jacobian is given by
$\mathbf{J}=(-d+u(\alpha-1))\mathbf{I}+u\tilde{\mathbf{A}}.$ (37)
$\lambda^{\prime}=-d+u(\alpha-1)+u\lambda^{\tilde{a}}.$ (38)
The bifurcation point of the attention-opinion bifurcation diagram occurs when the dominant eigenvalue of the Jacobian is zero, reaching the upper bound for stability of the equilibrium $\mathbf{x}_{e}$. As $u$ continues to increases and the dominant eigenvalue becomes positive, the equilibrium $\mathbf{x}_{e}$ become unstable and a bifurcation emerges. Solving for the critical value of the attention yields $u^{*}=\nicefrac{{d}}{{\left(\alpha-1+\lambda^{\tilde{a}}_{\max}\right)}}$. ∎
A.2.4 Lemma 5.5: BIMP converges to equilibrium